OpenStudy (anonymous):

Please help me write x^2+6x+8=0 in a vortex form ill give medal!!!!

OpenStudy (jdoe0001):

x^2+6x+8=0 written in a vortex will be x ^2 + 6x +8 = 0 - -- v vortex here sucking up the equation

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

Yes I believe so but if you could just re explain that to me that would be great

OpenStudy (jdoe0001):

\(\bf (a\pm b)^2\implies a^2\pm {\color{red}{ 2}}ab+b^2\) notice the middle term, is just 2 * "a" and "b" square rooted

OpenStudy (anonymous):

oh ok

OpenStudy (jdoe0001):

so let's grab yours and do some grouping first \(\bf x^2+6x+8=0\implies (x^2+6x)+8=0\implies (x^2+6x+{\color{red}{ \square }}^2)+8=0\) so what do you think we need there to get a "perfect square trinomial" ?

OpenStudy (anonymous):

b?

OpenStudy (jdoe0001):

let's see that \(\large \begin{array}{llll} (x^2+&6x+{\color{red}{ b }}^2)\\ &\uparrow \\ &2\cdot x\cdot b\ne 6x \end{array}\)

OpenStudy (jdoe0001):

so is not "b"... so what do you think might be?

OpenStudy (anonymous):

is it a letter or number that's supposed to go there?

OpenStudy (jdoe0001):

well, the middle term is \(\bf 2\cdot \sqrt{\textit{term on its left}}\cdot \sqrt{\textit{term on its right}}\) whatever those tems might be I just happen to use "a" and "b", but they could be anything

OpenStudy (anonymous):

so it would be "a" your saying

OpenStudy (jdoe0001):

if you look at your equation, we can tell from the grouping that the middle term is 6x so 2 * "term on the left" * "term on the right" has to equal 6x

OpenStudy (jdoe0001):

yes, "a" and "b" just means, the left-term and right-term

OpenStudy (anonymous):

sooooo....... that was what you wanted right, a, because honestly I don't know what it is you want from me now

OpenStudy (jdoe0001):

we need to find a value for the right-side-term

OpenStudy (anonymous):

What is the value supposed to be?

OpenStudy (anonymous):

I don't understand:(

OpenStudy (jdoe0001):

hmm the system is getting lagged

OpenStudy (anonymous):

it is actually because your user want showing

OpenStudy (jdoe0001):

lemme rewrite it a bit \(\bf 2\cdot x\cdot {\color{red}{ \square }}=6x\) if you solve for \(\bf {\color{red}{ \square }}\) what would you get?

OpenStudy (anonymous):

8?

OpenStudy (jdoe0001):

let's try 8 2 * x * 8 = 6 * x 16x \(\ne\) 6x

OpenStudy (anonymous):

so its wrong?

OpenStudy (jdoe0001):

well, 16x is not equal 6x..... so is not 8 have you covered simplifying linear equations yet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its 3

OpenStudy (jdoe0001):

look at it this way 2 * a * b = 6ab <---- what would you get for "b"?

OpenStudy (jdoe0001):

right is 3 that means 2 * x * 3 = 6x <--- that IS your middle term in the trinomial \(\bf 2\cdot \sqrt{\textit{term on its left}}\cdot \sqrt{\textit{term on its right}}\)

OpenStudy (anonymous):

then the other side is 2

OpenStudy (jdoe0001):

so now we know what is the number we need in order to complete the perfect square trinomial now keep in mind that all we're doing is borrowing from "0", zero that is, if we ADD \(3^2\) we have to also SUBTRACT \(3^2\) so let us use 3 \(\bf x^2+6x+8=0\implies (x^2+6x)+8=0 \\ \quad \\ \implies (x^2+6x+{\color{red}{ 3 }}^2)+8-{\color{red}{ 3 }}^2=0\implies (x+3)^2+8-{\color{red}{ 3 }}^2=0\)

OpenStudy (jdoe0001):

8 - 9 = -1 so \(\bf (x+3)^2-9=0\) <--- would be the vertex form \(\bf (x+{\color{red}{ 3}})^2-{\color{blue}{ 9}}=0\implies (x-{\color{red}{ -3}})^2-{\color{blue}{ 9}}=0\qquad vertex\implies ({\color{red}{ -3}},{\color{blue}{ 9}})\)

OpenStudy (jdoe0001):

wooops, lemme fix that... darn =(

OpenStudy (jdoe0001):

8 - 9 = -1 \(\bf (x+{\color{red}{ 3}})^2-{\color{blue}{ 1}}=0\implies (x-{\color{red}{ -3}})^2-{\color{blue}{ 1}}=0\qquad vertex\implies ({\color{red}{ -3}},{\color{blue}{ 1}})\)

OpenStudy (jdoe0001):

hmm shoud be -1 shoot anyhow \(\bf (x+{\color{red}{ 3}})^2-{\color{blue}{ 1}}=0\implies (x-{\color{red}{ -3}})^2-{\color{blue}{ 1}}=0\qquad vertex\implies ({\color{red}{ -3}},{\color{blue}{ -1}})\)

OpenStudy (anonymous):

so the vertex form is (x+3)^2 -1=0

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thank u my friend and just as promised...............

OpenStudy (jdoe0001):

yw